# A Causal Perspective on Measuring, Explaining and Mitigating Smells in LLM-Generated Code

### v3

@@ Abstract.
journalyear: 2026copyright: ccconference: 2026 IEEE/ACM 48th International Conference on Software Engineering; April 12–18, 2026; Rio de Janeiro, Brazilbooktitle: 2026 IEEE/ACM 48th International Conference on Software Engineering (ICSE ’26), April 12–18, 2026, Rio de Janeiro, Brazildoi: 10.1145/3744916.3773164isbn: 979-8-4007-2025-3/2026/04ccs: Software and its engineering Software verification and validationccs: Computing methodologies Knowledge representation and reasoningccs: Computing methodologies Probabilistic reasoningccs: Theory of computation Semantics and reasoning
@@ 1. Introduction
LLMs have rapidly moved from experimental tools to everyday assistants in software engineering (SE) (watson2021systematicliteraturereviewuse). Developers rely on LLM-based tools for many tasks such as code completion (9616462; lin2024soen101codegenerationemulating; 7180092; sun2025bridgingdeveloperinstructionscode), summarization (ahmed2022fewshottrainingllmsprojectspecific; 11029737; 10.1145/3728949), program repair (jin2023inferfixendtoendprogramrepair; 10.1145/3650212.3680323), clone detection (10.1145/2970276.2970326; 10556419), and test generation (Watson_2020; 10329992; 10765033). As generated code increasingly enters production systems, evaluation must look beyond functional correctness and consider whether the output meets broader standards of maintainable and robust software.
Code smells provide a foundation for evaluating broader quality concerns by capturing recurring design and implementation issues that affect readability, complexity, and long-term maintainability (smells_michele; palomba2018diffuseness; Evtikhiev_2023). Prior work has traditionally attributed smell introduction to human-driven factors such as developer experience, time pressure, and project constraints (smells_michele). The growing involvement of LLMs challenges this assumption. Models trained on large-scale code corpora inherit patterns present in the training data, including smelly or suboptimal code (10795572; tambon2024bugslargelanguagemodels), and may introduce additional defects or vulnerabilities during generation (10.1145/3605098.3636058; pearce2021asleepkeyboardassessingsecurity). Smell formation now reflects a combination of human design choices and model behavior, which renders traditional explanations incomplete.
Current evaluation practices are not designed to reflect model decisions influencing smell generation. For instance, similarity metrics such as BLEU (papineni2002bleu) and CodeBLEU (codebleu), along with execution-based measures, offer useful information about functional correctness but provide limited insight into structural or design quality (empirical_code_smells; 10555682; nunes2025evaluatingeffectivenessllmsfixing; takerngsaksiri2025codereadabilityagelarge). Although static analysis tools (SATs) can identify smells in model outputs, post-hoc detection alone falls short in revealing why a model introduced smells or which components of the generation process shaped them. As a result, developers lack visibility into factors that influence the maintainability of LLM-generated code.
@@
While these questions remain open, recent work has begun to introduce early indicators of smell propensity in LLMs. Velasco et al. proposed the PSC metric, a probabilistic score that estimates the likelihood that a model generates tokens associated with a given smell (velasco2025propense). Introduced as part of a benchmarking effort, PSC provides a continuous signal of model tendencies and a promising foundation for investigating the factors behind smell occurrence. Yet the broader dynamics of smell generation remain unexplored. 1) The drivers of variation in PSC and their relationship to the structural quality of generated code are still unknown, and 2) the stability of the metric across diverse conditions remains untested. Finally, 3) PSC has yet to be used to uncover mechanisms that explain why models introduce smells or to support interventions that reduce smell propensity. These three gaps hinder a full understanding of how smells emerge in LLM-generated code and limit our ability to devise approaches for reducing their occurrence.
@@
  • We release a dataset, benchmark, and implementation to support future research on explaining and mitigating smell generation (psc_repo).
@@ 2. Motivation & Related Work
Limitations of current evaluations. Existing evaluation methods prioritize surface-level correctness or lexical similarity. As a result, outputs that pass canonical metrics may still exhibit design flaws or maintenance issues. Recent empirical studies have highlighted this gap. Siddiq et al. (empirical_code_smells) and Torek et al. (melin2024precisionperilevaluatingcode) showed that both the training data and the code generated from LLMs frequently exhibit code smells and insecure constructs. Alif et al. (mohsin2024trustlargelanguagemodels) observed that models like Bard and ChatGPT introduce more smells than Copilot or CodeWhisperer in complex tasks. Similarly, Kharma et al. (kharma2025securityqualityllmgeneratedcode) linked outdated training corpora to the persistence of legacy antipatterns. In the context of test generation, Ouédraogo et al. (ouedraogo2024testsmellsllmgeneratedunit) found high rates of test-specific smells in the outputs of GPT-3.5 and GPT-4. Although these findings underscore the need for deeper evaluation, current tools do not provide standardized methods to measure or explain the structural quality of LLM-generated code.
Toward principled evaluation and understanding. To address these limitations, recent work has introduced specialized benchmarks and metrics aimed at evaluating and explaining the structural quality of LLM-generated code. The CodeSmellEval benchmark (velasco2025propense) proposed the Propensity Smelly Score (PSC), a probabilistic metric that estimates the likelihood of generating specific types of smell. Zheng et al. (zheng2024correctnessbenchmarkingmultidimensionalcode) developed the RACE framework to incorporate broader quality dimensions such as maintainability and readability. Wu et al. (wu2024ismell) introduced iSMELL, a system that combines static analysis with LLMs to improve smell detection and refactoring. Although these approaches mark important progress, key questions remain unanswered: Which generation factors most influence the presence of smell? Can we isolate causal contributors to quality degradation? And can smell generation be systematically reduced through actionable strategies?
@@ 3. Propensity Smelly Score (PSC)
The Propensity Smelly Score (PSC) is a probabilistic metric that estimates the likelihood that a LLM generates a specific type of code smell (velasco2025propense). It relies on the next-token probabilities predicted by the model during autoregressive generation. Given a token sequence $w=\{w_{1},w_{2},\ldots,w_{n}\}$, the model produces a probability distribution in the vocabulary for each token position. The PSC computation process is depicted in Fig. 1.
@@
The function $\theta_{\mu}$ reflects the model’s average confidence in generating the sequence associated with a given code smell. Higher PSC values indicate a stronger tendency to produce tokens that correspond to known smells. To support comparisons across models, we adopt a threshold $\lambda=0.5$ from prior work (velasco2025propense) as a default interpretability criterion, where a PSC value at or above $\lambda$ indicates that the model is likely to generate the corresponding smell. This threshold is used only to interpret the score and does not affect how PSC is computed. Adjusting it changes how many instances are labeled as propense but leaves the underlying probabilities unchanged.
@@ 4.1. Semantic Preserving Code Transformations
Our methodology draws on prior work in robustness evaluation for neural program repair using natural transformations (sept_paper). We focus specifically on statement-level transformations. Broader edits, such as converting while loops to for loops (While2For), can significantly alter the code’s structural context, thereby influencing the detection of code smells. Limiting the analysis to statement-level changes preserves semantics and overall structure while still introducing meaningful syntactic variation.
@@ 4.2. Information Gain (IG)
BLEU and CodeBLEU are widely used in code generation benchmarks and are often treated as default baselines. Both rely on surface-level similarity, using n-gram overlap and syntax-based matching to assess how closely a generated snippet aligns with a reference. These metrics are effective for evaluating correctness but are not designed to reflect deeper structural properties such as code smells. Because no established alternatives exist for assessing structural quality in generated code, we include BLEU and CodeBLEU as representative standards for comparison (Evtikhiev_2023; melin2024precisionperilevaluatingcode). Their reliance on reference similarity contrasts with PSC, which captures likelihood-based signals from the model itself. To compare metrics that draw on such different sources of information, we adopt an information-gain framework that quantifies how much each metric reduces uncertainty about the severity of code smells.
@@ 4.3. Causal Analysis of Factors Influencing PSC
Previous studies support the inclusion of these variables. Sampling strategies with high temperature tend to produce longer and more deeply nested code, whereas greedy decoding yields shorter and flatter sequences (donato2025studyingconfigurationsimpactcode). Model size shapes output as well. Larger models typically produce more concise and focused completions, while smaller ones often generate repetitive or verbose code (chen2021evaluatinglargelanguagemodels). Architectural design introduces further variation. Models that use grouped-query attention and sliding windows, such as Mistral, often produce shorter sequences than similarly sized LLaMA models (jiang2023mistral7b). Prompt formulation has a comparable effect. Detailed prompts, including full function signatures, tend to elicit longer and more structured outputs, whereas vague prompts result in shorter completions (khojah2024impactpromptprogrammingfunctionlevel). Finally, probabilistic metrics reflect the underlying structure of code. The ASTTrust study (palacio2024syntax) shows that metrics such as PSC vary with AST complexity and token alignment.
@@ 4.3.1. Computing Average Treatment Effects
The average treatment effect (ATE), shown in Fig. 2, captures the expected change in PSC resulting from an intervention, averaged over the distribution of confounders. Formally, ATE is expressed as $P(Y\mid\text{do}(T))$ in Pearl’s do-calculus (Pearl2018Causality). We identify valid causal estimands by applying graphical criteria (e.g., back-door criterion), which determines the appropriate adjustment set $Z$ from the causal graph.
@@ 4.5.2. Sampling
We followed a purpose sampling strategy (sampling), since our objective was to collect early evidence and practitioner feedback on the usefulness of PSC to assess the propensity to generate code smells at the snippet level. The participants were selected based on their combined experience in machine learning and software engineering. We prioritized relevant expertise over randomness to ensure a meaningful interpretation of the metric output. Invitations were sent by email across academic and professional networks, and recipients were encouraged to share the survey with peers.
@@ 4.5.5. Statistical Analysis
To evaluate the impact of PSC exposure on participant responses, we applied non-parametric statistical tests suitable for ordinal data. Specifically, we used the Mann–Whitney U test to compare responses between the control and treatment groups for each of the three main questions ($Q_{1}$$Q_{3}$). This test does not assume normality and is appropriate for detecting differences in central tendency between independent samples with Likert scale results. Statistical comparisons were performed independently for each question and code smell type, yielding a total of 15 comparisons (3 questions in 5 snippets). For each comparison, we report the $p$ value and highlight statistically significant differences ($p<.05$). Open-ended responses from $Q_{5}$ were analyzed using thematic analysis (themes), with two authors independently coding answers and consolidating recurring ideas into broader themes.
@@ 5. Experimental Context
$M_{1}$ CodeLlama-7b-hf (codeLlama) 7B $S_{1}$ Qwen2.5-Coder-0.5B (qwen) 0.5B
$M_{2}$ Mistral-7B-v0.3 (mistral) 7B $S_{2}$ Qwen2.5-Coder-1.5B (qwen) 1.5B
$M_{3}$ Qwen2.5-Coder-7B (qwen) 7B $S_{3}$ Qwen2.5-Coder-3B (qwen) 3B
$M_{4}$ starcoder2-7b (starcoder) 7B $S_{4}$ Qwen2.5-Coder-7B (qwen) 7B
@@
InfixDividing
Divide an in/pre/post-fix
expression into two expressions
x = a + b * c
temp = b *c
x = a + temp
@@ 5.1. Dataset
Our experiments are based on an expanded version of the CodeSmellData corpus (velasco2025propense), referred to as CodeSmellData 2.0. This dataset was constructed through a multistage pipeline designed to extract and annotate large-scale Python code snippets from open-source repositories, allowing empirical evaluation of code smell generation in LLM-generated code.
Data Mining and Processing. We began by extracting method-level Python snippets from public GitHub repositories with more than $1K$ stars, published between January 2020 and January 2025. This process yielded approximately $4.7M$ methods, which were stored in a MySQL database. Each snippet was processed to compute syntactic and lexical confounders using the GALERAS schema (galeras). In addition to these features, we extracted part-of-speech (POS) token counts using the spaCy parser to support fine-grained structural and linguistic analysis.
Deduplication and Annotation. We removed exact and semantically similar duplicates by comparing abstract syntax tree (AST) representations, resulting in a deduplicated set of approximately $2.1M$ unique methods exported to JSON format. Methods that did not perform the AST parsing were excluded. The remaining snippets were analyzed using Pylint to identify code smells, recording their types and sources. This step produced around $2.8M$ unique code smell instances, which collectively comprise CodeSmellData 2.0 (psc_repo).
@@
Table 3 lists the included smell types and their Pylint identifiers. Each type is capped at 500 instances for consistency. However, the full distribution of smells from CodeSmellData 2.0 is preserved in the dataset released (psc_repo).
@@ 5.3. Causal Settings
The $T_{1}$ intervention (see 1 in Fig. 3) manipulates the generation strategy used by the language model during inference. We evaluated six decoding methods: greedy search, beam search, sampling, contrastive search, top-$k$ sampling, and top-$p$ sampling. The beam search is configured with five beams and early stopping enabled. Contrastive search uses a penalty of 0.6 and a top-$k$ value of 4. The topkk sampling is applied with $k=50$, and the toppp sampling uses $p=0.9$. All decoding strategies are applied to the same model instance ($M_{1}$) to ensure comparability between conditions.

### v4

@@ Abstract.
journalyear: 2026copyright: ccconference: 2026 IEEE/ACM 48th International Conference on Software Engineering; April 12–18, 2026; Rio de Janeiro, Brazilbooktitle: 2026 IEEE/ACM 48th International Conference on Software Engineering (ICSE ’26), April 12–18, 2026, Rio de Janeiro, Brazildoi: 10.1145/3744916.3773164isbn: 979-8-4007-2025-3/2026/04submissionid: icse2026-p2419ccs: Software and its engineering Software verification and validationccs: Computing methodologies Knowledge representation and reasoningccs: Computing methodologies Probabilistic reasoningccs: Theory of computation Semantics and reasoning
@@ 1. Introduction
LLMs have rapidly moved from experimental tools to everyday assistants in software engineering (SE) (Watson et al., 2021). Developers rely on LLM-based tools for many tasks such as code completion (Ciniselli et al., 2022; Lin et al., 2024; White et al., 2015; Sun et al., 2025b), summarization (Ahmed and Devanbu, 2022; Sun et al., 2025a; Fang et al., 2025), program repair (Jin et al., 2023; Xia and Zhang, 2024), clone detection (White et al., 2016; Khajezade et al., 2024), and test generation (Watson et al., 2020; Schäfer et al., 2024; Yang et al., 2024). As generated code increasingly enters production systems, evaluation must look beyond functional correctness and consider whether the output meets broader standards of maintainable and robust software.
Code smells provide a foundation for evaluating broader quality concerns by capturing recurring design and implementation issues that affect readability, complexity, and long-term maintainability (Tufano et al., 2015; Palomba et al., 2018; Evtikhiev et al., 2023). Prior work has traditionally attributed smell introduction to human-driven factors such as developer experience, time pressure, and project constraints (Tufano et al., 2015). The growing involvement of LLMs challenges this assumption. Models trained on large-scale code corpora inherit patterns present in the training data, including smelly or suboptimal code (Ramírez et al., 2024; Tambon et al., 2024), and may introduce additional defects or vulnerabilities during generation (Aljohani and Do, 2024; Pearce et al., 2021). Smell formation now reflects a combination of human design choices and model behavior, which renders traditional explanations incomplete.
Current evaluation practices are not designed to reflect model decisions influencing smell generation. For instance, similarity metrics such as BLEU (Papineni et al., 2002) and CodeBLEU (Ren et al., 2020), along with execution-based measures, offer useful information about functional correctness but provide limited insight into structural or design quality (Siddiq et al., 2022, 2024; Nunes et al., 2025; Takerngsaksiri et al., 2025). Although static analysis tools (SATs) can identify smells in model outputs, post-hoc detection alone falls short in revealing why a model introduced smells or which components of the generation process shaped them. As a result, developers lack visibility into factors that influence the maintainability of LLM-generated code.
@@
While these questions remain open, recent work has begun to introduce early indicators of smell propensity in LLMs. Velasco et al. proposed the PSC metric, a probabilistic score that estimates the likelihood that a model generates tokens associated with a given smell (Velasco et al., 2025). Introduced as part of a benchmarking effort, PSC provides a continuous signal of model tendencies and a promising foundation for investigating the factors behind smell occurrence. Yet the broader dynamics of smell generation remain unexplored. 1) The drivers of variation in PSC and their relationship to the structural quality of generated code are still unknown, and 2) the stability of the metric across diverse conditions remains untested. Finally, 3) PSC has yet to be used to uncover mechanisms that explain why models introduce smells or to support interventions that reduce smell propensity. These three gaps hinder a full understanding of how smells emerge in LLM-generated code and limit our ability to devise approaches for reducing their occurrence.
@@
  • We release a dataset, benchmark, and implementation to support future research on explaining and mitigating smell generation (Lab, 2024).
@@ 2. Motivation & Related Work
Limitations of current evaluations. Existing evaluation methods prioritize surface-level correctness or lexical similarity. As a result, outputs that pass canonical metrics may still exhibit design flaws or maintenance issues. Recent empirical studies have highlighted this gap. Siddiq et al. (Siddiq et al., 2022) and Torek et al. (Melin et al., 2024) showed that both the training data and the code generated from LLMs frequently exhibit code smells and insecure constructs. Alif et al. (Mohsin et al., 2024) observed that models like Bard and ChatGPT introduce more smells than Copilot or CodeWhisperer in complex tasks. Similarly, Kharma et al. (Kharma et al., 2025) linked outdated training corpora to the persistence of legacy antipatterns. In the context of test generation, Ouédraogo et al. (Ouédraogo et al., 2024) found high rates of test-specific smells in the outputs of GPT-3.5 and GPT-4. Although these findings underscore the need for deeper evaluation, current tools do not provide standardized methods to measure or explain the structural quality of LLM-generated code.
Toward principled evaluation and understanding. To address these limitations, recent work has introduced specialized benchmarks and metrics aimed at evaluating and explaining the structural quality of LLM-generated code. The CodeSmellEval benchmark (Velasco et al., 2025) proposed the Propensity Smelly Score (PSC), a probabilistic metric that estimates the likelihood of generating specific types of smell. Zheng et al. (Zheng et al., 2024) developed the RACE framework to incorporate broader quality dimensions such as maintainability and readability. Wu et al. (Wu et al., 2024) introduced iSMELL, a system that combines static analysis with LLMs to improve smell detection and refactoring. Although these approaches mark important progress, key questions remain unanswered: Which generation factors most influence the presence of smell? Can we isolate causal contributors to quality degradation? And can smell generation be systematically reduced through actionable strategies?
@@ 3. Propensity Smelly Score (PSC)
The Propensity Smelly Score (PSC) is a probabilistic metric that estimates the likelihood that a LLM generates a specific type of code smell (Velasco et al., 2025). It relies on the next-token probabilities predicted by the model during autoregressive generation. Given a token sequence $w=\{w_{1},w_{2},\ldots,w_{n}\}$, the model produces a probability distribution in the vocabulary for each token position. The PSC computation process is depicted in Fig. 1.
@@
The function $\theta_{\mu}$ reflects the model’s average confidence in generating the sequence associated with a given code smell. Higher PSC values indicate a stronger tendency to produce tokens that correspond to known smells. To support comparisons across models, we adopt a threshold $\lambda=0.5$ from prior work (Velasco et al., 2025) as a default interpretability criterion, where a PSC value at or above $\lambda$ indicates that the model is likely to generate the corresponding smell. This threshold is used only to interpret the score and does not affect how PSC is computed. Adjusting it changes how many instances are labeled as propense but leaves the underlying probabilities unchanged.
@@ 4.1. Semantic Preserving Code Transformations
Our methodology draws on prior work in robustness evaluation for neural program repair using natural transformations (Le-Cong et al., 2024). We focus specifically on statement-level transformations. Broader edits, such as converting while loops to for loops (While2For), can significantly alter the code’s structural context, thereby influencing the detection of code smells. Limiting the analysis to statement-level changes preserves semantics and overall structure while still introducing meaningful syntactic variation.
@@ 4.2. Information Gain (IG)
BLEU and CodeBLEU are widely used in code generation benchmarks and are often treated as default baselines. Both rely on surface-level similarity, using n-gram overlap and syntax-based matching to assess how closely a generated snippet aligns with a reference. These metrics are effective for evaluating correctness but are not designed to reflect deeper structural properties such as code smells. Because no established alternatives exist for assessing structural quality in generated code, we include BLEU and CodeBLEU as representative standards for comparison (Evtikhiev et al., 2023; Melin et al., 2024). Their reliance on reference similarity contrasts with PSC, which captures likelihood-based signals from the model itself. To compare metrics that draw on such different sources of information, we adopt an information-gain framework that quantifies how much each metric reduces uncertainty about the severity of code smells.
@@ 4.3. Causal Analysis of Factors Influencing PSC
Previous studies support the inclusion of these variables. Sampling strategies with high temperature tend to produce longer and more deeply nested code, whereas greedy decoding yields shorter and flatter sequences (Donato et al., 2025). Model size shapes output as well. Larger models typically produce more concise and focused completions, while smaller ones often generate repetitive or verbose code (Chen et al., 2021). Architectural design introduces further variation. Models that use grouped-query attention and sliding windows, such as Mistral, often produce shorter sequences than similarly sized LLaMA models (Jiang et al., 2023a). Prompt formulation has a comparable effect. Detailed prompts, including full function signatures, tend to elicit longer and more structured outputs, whereas vague prompts result in shorter completions (Khojah et al., 2024). Finally, probabilistic metrics reflect the underlying structure of code. The ASTTrust study (Palacio et al., 2024) shows that metrics such as PSC vary with AST complexity and token alignment.
@@ 4.3.1. Computing Average Treatment Effects
The average treatment effect (ATE), shown in Fig. 2, captures the expected change in PSC resulting from an intervention, averaged over the distribution of confounders. Formally, ATE is expressed as $P(Y\mid\text{do}(T))$ in Pearl’s do-calculus (Pearl and Mackenzie, 2018). We identify valid causal estimands by applying graphical criteria (e.g., back-door criterion), which determines the appropriate adjustment set $Z$ from the causal graph.
@@ 4.5.2. Sampling
We followed a purpose sampling strategy (Baltes and Ralph, 2021), since our objective was to collect early evidence and practitioner feedback on the usefulness of PSC to assess the propensity to generate code smells at the snippet level. The participants were selected based on their combined experience in machine learning and software engineering. We prioritized relevant expertise over randomness to ensure a meaningful interpretation of the metric output. Invitations were sent by email across academic and professional networks, and recipients were encouraged to share the survey with peers.
@@ 4.5.5. Statistical Analysis
To evaluate the impact of PSC exposure on participant responses, we applied non-parametric statistical tests suitable for ordinal data. Specifically, we used the Mann–Whitney U test to compare responses between the control and treatment groups for each of the three main questions ($Q_{1}$$Q_{3}$). This test does not assume normality and is appropriate for detecting differences in central tendency between independent samples with Likert scale results. Statistical comparisons were performed independently for each question and code smell type, yielding a total of 15 comparisons (3 questions in 5 snippets). For each comparison, we report the $p$ value and highlight statistically significant differences ($p<.05$). Open-ended responses from $Q_{5}$ were analyzed using thematic analysis (Cruzes and Dyba, 2011), with two authors independently coding answers and consolidating recurring ideas into broader themes.
@@ 5. Experimental Context
$M_{1}$ CodeLlama-7b-hf (Rozière et al., 2024) 7B $S_{1}$ Qwen2.5-Coder-0.5B (Qwen et al., 2025) 0.5B
$M_{2}$ Mistral-7B-v0.3 (Jiang et al., 2023b) 7B $S_{2}$ Qwen2.5-Coder-1.5B (Qwen et al., 2025) 1.5B
$M_{3}$ Qwen2.5-Coder-7B (Qwen et al., 2025) 7B $S_{3}$ Qwen2.5-Coder-3B (Qwen et al., 2025) 3B
$M_{4}$ starcoder2-7b (Lozhkov et al., 2024) 7B $S_{4}$ Qwen2.5-Coder-7B (Qwen et al., 2025) 7B
@@
InfixDividing Divide an in/pre/post-fix expression into two expressions x = a + b * c temp = b *c x = a + temp
@@ 5.1. Dataset
Our experiments are based on an expanded version of the CodeSmellData corpus (Velasco et al., 2025), referred to as CodeSmellData 2.0. This dataset was constructed through a multistage pipeline designed to extract and annotate large-scale Python code snippets from open-source repositories, allowing empirical evaluation of code smell generation in LLM-generated code.
Data Mining and Processing. We began by extracting method-level Python snippets from public GitHub repositories with more than $1K$ stars, published between January 2020 and January 2025. This process yielded approximately $4.7M$ methods, which were stored in a MySQL database. Each snippet was processed to compute syntactic and lexical confounders using the GALERAS schema (Rodriguez-Cardenas et al., 2023). In addition to these features, we extracted part-of-speech (POS) token counts using the spaCy parser to support fine-grained structural and linguistic analysis.
Deduplication and Annotation. We removed exact and semantically similar duplicates by comparing abstract syntax tree (AST) representations, resulting in a deduplicated set of approximately $2.1M$ unique methods exported to JSON format. Methods that did not perform the AST parsing were excluded. The remaining snippets were analyzed using Pylint to identify code smells, recording their types and sources. This step produced around $2.8M$ unique code smell instances, which collectively comprise CodeSmellData 2.0 (Lab, 2024).
@@
Table 3 lists the included smell types and their Pylint identifiers. Each type is capped at 500 instances for consistency. However, the full distribution of smells from CodeSmellData 2.0 is preserved in the dataset released (Lab, 2024).
@@ 5.3. Causal Settings
[…]